OpenStudy (anonymous):

the equation of the line tangent to the curve y= kx+8/k+x at x=-2 is y=x+4 what is the value of k?

OpenStudy (anonymous):

do you know the answer?

OpenStudy (anonymous):

i feel like its 1

OpenStudy (anonymous):

okay, i just check and i know now, we have an equation where the k does not play into account, that is the equation of the tangent, what do we know of the tangent? is that at point(x,y) which in this case (-2,y) we have an tangent for y=kx+8/k+x so using the equation of the tangent we can work out y which is y=-2+4 = 2 so we have (-2,2) now, we have two knowns and 1 unknown in the original equation of the curve, if we plug these in we get y=(-2k+8)/(k-2)=2 we know that x=-2 ==> -2k+8=2k-4 ==> 12=4k ==> k=3

OpenStudy (anonymous):

do you understand?

OpenStudy (anonymous):

look at picture, red line is the line of the tangent where they touch is (-2,2) the blue line is the original equation but with k=3, so this is correct

OpenStudy (anonymous):

yes thank you! i dropped a negative sign so it didnt make sense! thank you your explenation was very helpfull!

OpenStudy (anonymous):

yw